PS/SPS User Meeting
Minutes of Meeting
The template for the minutes of meeting for the weekly PS & SPS users meeting can always be found using this link: https://codimd.web.cern.ch/0aHs5uqmSPuII76-SaaQdg
Please pre-fill them before the meeting with a summery of your activities!
PS-SPS Users Meeting for Week 46 held on November 14th 2024
Coordinator for this week: E. Barbara Holzer
- Submission period for beam requests for 2025 did close on October 21st, please get in touch urgently if you missed the deadline or still have open questions.
-
Updated user schedule v4.1.0 on October 28th
-
Please be aware that the infrastructure work (installations/de-installations) for the FOLLOWING week are FINALIZED by the EA-team the Thursday early afternoon in the week BEFORE
-
DRAFT Injector schedule for 2025 is available here: https://edms.cern.ch/document/3057281/latest
-
AD/ELENA is installing a new beam line for antiproton / H- test beam called TELMAX. Call for beam requests for the TELMAX beamline for 2024 and 2025 is expected to go out soon.
EuroLabs : please urgently submit a list of your publications to Tanya if you profited from EuroLabs funding.
News from the PS & SPS Physics Coordinator
E.B. Holzer, M.R. Jäkel
- 2024 injector schedule released EDMS 2872566
2024 approved LHC schedule EDMS 2872429
Proton Run 2024
- AD/ELENA Physics Stop Monday 02.12.2024
- SPS NA Physics Stop protons Thursday 31.10.2024
- PS EA Physics Stop protons Wednesday 27.11.2024
Ion Run 2024
- SPS NA Physics Start 4.11.2024
Stop Pb ions Monday 2.12.2024 (6h)- NA : Week 45-47 high energy
- NA : Week 48 low energy (no LHC running)
- PS EA Physics Start 6.11.2024 (tbd)
Stop Pb ions Monday 2.12.2024 (6h)- CHIMERA: 13.11-2.12
Reminder: Beam time exceeding the limits of 2 weeks PS beam time and 1 week SPS beam time per year need the approval of one of these CERN committees: SPSC, LHCC, DRDC, INTC, RB or IEFC. Consider joining a DRD collaboration, if you require more beam time.
Target intensities:
See below or via ASM (accessible from within the CERN network).
News from the Facilities Operations Meeting (FOM)
Upcoming SPS MDs next weeks :
https://be-dep-ea.web.cern.ch/content/md-planning-north-area
Dedicated SPS MDs - if not anounced differently - are taking place from 8:00 to 18:00
- Week 45: No dedicated MD
- Week 46: Ions Dedicated MD (Wed)
- Week 47: Ions Dedicated MD (Wed) & Long parallel (Thu), moved from week 46
PS Machine Report (R. Garcia Alia)
- 96.1% Thu-to-Thu availability
- No doubleTOF programmed to TOF anymore (due to target safety considerations - to be followed up)
- Recent check of T8 beam stopper
- End of protons and start of ions in EAST T8
- Some LEIR issues for NOMINAL beam on Wed night - E-cooler trimmed to adapt to lower intensity
SPS Machine Report (T. Levens)
- A relatively good week with 95% availability for NA
- Few faults, but long periods of LHC filling with no beam to NA
- Current schedule for LHC MD block:
- Today (until 23h)
- Tomorrow afternoon (11h to 18h)
- Monday 18th afternoon (16h to 19h)
- Sunday 24th morning until Monday 25th morning
- 4h LHC MD preparation in SPS planned on Friday 22nd (morning)
- Beam to NA will be provided whenever possible outside of LHC filling
- Today’s SPS long parallel MD has been delayed to next week
Safety (A. Schouten)
ISIEC Safety Clearence (please login to Drupal): https://ep-th-safety.web.cern.ch/isiec-safety-clearance
- General reminder PPE compliance. ISIEC requests have been scheduled in and will be handled by the person who responded to the request.
n_ToF (M.Bacak, A.Manna, P.M.Milazzo)
Smooth data taking in all experimental areas during week #45
- EAR1: 65Cu(n, g) with C6D6 detectors
- EAR2: (n, lcp) studies
- Activation measurement in the NEAR
On 11.11 change of neutron collimator in EAR2 from the one for capture to a larger one for fission.
Mounting of a a FIssion Fragments Identification arm (FIFI).
FIFI can measure mass and energy of emitted fragments by means of a couple of microchannel plates and an ionization chamber.
Week #46
EAR1: C6D6 detectors to measure the 65Cu(n, g) reaction
EAR2: fission fragments study
Both measurements will run up to end of the 2024 campaign (on 25.11)
Received protons
AD: PAX (M. Roosa)
Preliminary data analysis from the test beam that concluded on November 8th suggests that the observed peaks in the HPGe detector are coming from pion activation of the Ge crystal itself. Sharp peak at 54 keV is likely a low-lying excited state in 73Ge.
Will conduct another short test the week of November 18th to further investigate this hypothesis and try to reduce further the instantaneous antiproton intensity to see the desired antiprotonic atom x rays.
East Area Beam Status (L. Nevay)
On-call number for beam physicists: 67500
T09/T10/T11: Good operation.
East Area Users Tour de Table
T8 Main: IRRAD/CHARM (F. Ravotti / S. Fiore)
Very good last week of the proton run (ended yesterday at 6am), see plot below. Despite long ion cycles for the LHC filling, both p/w and TID reference values reached. Also beam centering was very good on both axis (~96% of the spill within +/- 2mm from the beam center).
Reference weekly intensity reached basically every week. The high-intensity spill operational this year significantly increased the intensity delivered to T8 (average daily intensity doubled since the LS2) for the benefit of the users and irradiation experiments carried out of both our irradiation areas. Therefore, we would like to warmly THANK all teams (across several CERN departments) involved in our operation for the continuous help and support to our facility!
Statistics 2024 for proton and mix-field irradiations:
All irradiation experiments completed in IRRAD and CHARM Access on Wednesday concernend only IRRAD to remove last samples and finalize the preparation of IRRAD Zone 1 for HEARTS:
Ion beam commissioning started yesterday, users starting in the coming days. Next week report will be made by HEARTS team.
T9 Main: WCTE (M. Hartz)
We continue to collect and analyze data for our beam monitors. Working towards particle identification using data:
Continue to work on the synchronous readout of the detector so we can collect useful physics data. Calibration runs with flashed LEDs are being run, but these can be done simultaneously with reception of the beam. We continue to use the beam to work on the synchronous readout of the detectors and beam monitors. We expect ~1 per day access of about 1 hour to change beam equipment (aerogels for different beam momenta).
T10 Outgoing Main: SHiP LS-SBT (A. Hollnagel)
SHiP LS-SBT exposure successfully concluded (Wed. 2024-11-06 – Wed. 2024-11-13):
-
Pre-assembly at T10 construction zone already on Tue. 2024-11-05, swift setup in T10 on Wed. 2024-11-06 afterwards. (Many thanks to the transport group for DESY table management and transport of our equipment in the morning! Many thanks for the safety group for giving us clearance in the evening so we could start first measurements during the night!)
-
Data taking with muons. Beam diagnostics with Cherenkov detector and tuning with collimators at the start to control hadron beam contamination. (Many thanks to Maarten!)
-
Data taking quite slow due to LHC filling, but all planned measurements successfully conducted. :)
-
Swift disassembly on Wed. 2024-11-06. (Again many thanks to the transport group!)
![]() |
![]() |
![]() |
---|---|---|
setup at PS EA T10 | beam telescope | liquid handling system |
3x 1-cell full-size LS-SBT prototypes (2 on DESY tables & rotation platforms):
- 120cm x 80cm x 20cm
- ~200l of Liquid scintillator (LS) made from linear alkylbenzene (LAB) + 2,5 diphenyloxazole (PPO) per detector cell
- 2 Wavelength-shifting Optical Module (WOM) light sensors with SiPM readout per detector cell
- Comparison of detector performance (overall light yield, WOM light share) with different inner cell surface reflectivities:
- A: AlMg4.5 (polished)
- B: CRES + PTFE liner (0.5mm)
- C: AlMg4.5 (raw)
- Test of new readout electronics at the end
![]() |
![]() |
![]() |
![]() |
![]() |
---|---|---|---|---|
front | back | inner surface (Al4.5 polished) | inner surface (CRES + PTFE liner) | readout |
Beam diagnostics with Cherenkov detectors & beam telescope:
1.1 bar | 1.5 bar |
---|---|
![]() |
![]() |
black ~muons, blue ~pions, red ~protons/kaons | black ~muons, blue ~pions, red ~protons/kaons |
T10 Main: LHCb (Please Put Your Name Here)
T11 Main: CLOUD (Please Put Your Name Here)
North Area Beam Status (L. Nevay)
On-call number for beam physicists: 67500
Target sharing:
T2 | T4 |
---|---|
300 x 1088 | 300 x 1088 |
General: A leak was found in TCC2 last week and found to be on the XTAX in front of H8/H6. Towards the end of the week the rate increased and an inspection was done. Given the relatively low intensity of the ion beams we can run without cooling until the end of the run. The leak can be properly fixed (with lower dose) during the YETS. Beam came back on Friday evening.
H2: Good operation.
H4: Good operation.
H8: Good operation.
H6/P42/K12/M2: No operation.
North Area Users Tour de Table
H2:
Main: NA61 SHINE (B. Maksiak, A. Dmitriev)
- Our detector works and the run goes well as long as there is the beam.
- LHC fillings take a significant beam time. Their not accurate announcements bring frustration to us. The beginnings and endings of the fillings could be announced earlier and updated in case there are changes of the plan.
- The intensity of beam is was unstable last week. Our beam expert is waken up two-three times per night.
- An interesting case of periodicity of beam intensity
- The spill structure (far from flat as reported a week ago) improved significantly - thank you!
- We are missing the beam extration beeping on the beam monitor TV broadcast, we frequently miss first couple of minutes after supercycle changes
- When the Pb source refilling is planned?
- We plan have the beam momentum changed to 13.5 GeV/c in the morning of Monday, 25.11 and we would like to have 2-3 hours access to PPE112.
H4:
Main: RE44 HERD (Wen-xi,Peng)
We finished the beam test as the main user. Some new results are presented here.
-Charge spectrum of PSD(readout by SiPM+ASIC, Z=1~35),
-Charge spectrum of CPID(small SSD+ASIC, Z=2~54),
Thanks again for all the supports from CERN.
Thanks to our Italian friends(AMS-L0 team from INFN/UniPG, Charge tagger team from INFN/UniFI)for providing the ps trigger, high precise measurements of track and charge.
Main: RE1 AMS L0 (M. Duranti - INFN Perugia)
- we didn’t take too many data: LHC fill, MD, etc…
- why the start of ion beams is on Monday and not on Wednesday? A full day of potential good beam spent in installing and a full day of MD is quite frustrating… Then add that if you do not have the dosimeter this forces you to be here on Thursday evening (or to not be at NA, on Monday, before 09:00 - 09:30 or so)
- we had to Patrol the full line (PPE124 “locked”, PPE144 “free access”);
- the insentity of the beam was too high, and we also saw a certain change after each LHC fill. Now will try to find a good compromise between the AMS-L0 detectors (Si ustrip + Charge Tagger) and the VLAST colleagues;
- the beam references on the line are misleading: many different references (maybe users too pro-active?)
- last “complain” (sorry!): we had the installation of the beam pipe with a guy only speaking french. We’re italian and we survive, but in general is not easy in this way…
Parallel: VLAST (LI, Xiang)
We’ve taken some data (not too many due to LHC fillings/MDs) from the beam, some of the charges can be tagged.
The beam intensity was too high in previous days, which caused some issues for another charge tagger. But Matteo reduced the intensity last night. We’d check the performance via the data soon.
Incoming Main: RE25 CALET (Please Put Your Name Here)
H8:
Outgoing Main: MEDIPIX (H. Cintas)
The campaign was succesful. We were able to validate our new Timepix4 readout and our satellite radiation monitor, the Hardpix. We achieved all our secondary objectives, even with the complex week we had.
Thanks to Maarten and Johannes for all their help during the week.
Our setup was packed and cleared from RP. Thanks to RP for their reactivity on Monday.
Outgoing Parallel: PAN (Please Put Your Name Here)
Main: RE7 GAMMA MEV (Please Put Your Name Here)
Incoming Main: NA60+ (E. Scomparin)
The set-up will be quite similar to that of 2023, a box with a telescope of Si detector prototypes, positioned on the beam axis, and prototypes of MWPC positioned at a small angle (10-20 degrees). Two DESY tables were requested for the installation. A 5mm Pb target will be used to produce secondary particles.
The program (1 week at 150 AGeV and 1 week at low energy) includes: (i) test of the prototypes, (ii) measurements of the beam optics at the two energies up to the nominal beam intensity foreseen for the experiment (10^7 Pb/spill), under guide of Maarten and colleagues and RP group, (iii) a measurement of charged hadron multiplicity in Pb-A collisions at 13.5A GeV.
Parasitic Users
H4 RE44 HERD (Wen-Xi, Peng)
In Monday, we just dismounted the Calorimeter, and move the rest detectors to the downstream of AMS L0. In this week, we will continue the data acquisition mostly for studying the performce of the silicon strip detectors with different parameters setup.
AoB
Minutes by the respective speakers, edited by E. B. Holzer, M. Jäkel, T. Shulha, and M. Schwinzerl